Financial Statements of a Manufacturing Firm

The following events took place for Digital Vibe Manufacturing Company during January, the first month of its operations as a producer of digital video monitors:

  1. Purchased $54,600 of materials
  2. Used $42,000 of direct materials in production.
  3. Incurred $62,800 of direct labor wages.
  4. Incurred $88,500 of factory overhead.
  5. Transferred $146,900 of work in process to finished goods.
  6. Sold goods for $262,600.
  7. Sold goods with a cost of $116,800.
  8. Incurred $67,200 of selling expenses.
  9. Incurred $29,500 of administrative expense.

Using the information given, complete the following:

a. Prepare the January income statement for Digital Vibe Manufacturing Company.

Digital Vibe Manufacturing Company
Income Statement
For the Month Ended January 31
Revenues $
Cost of goods sold
Gross profit $
Operating expenses:
Selling expenses $
Administrative expenses
Total operating expenses
Net income $

Feedback

b. Determine the Materials Inventory, Work in Process Inventory, and Finished Goods Inventory balances at the end of the first month of operations.

Digital Vibe Manufacturing Company
Inventory Balances
For the Month Ended January 31
Inventory balances on January 31:
Materials $
Work in process
Finished goods
Answer #1
a
Digital Vibe Manufacturing Company
Income Statement
For the Month Ended January 31
Revenues 262600
Cost of goods sold 116800
Gross profit 145800
Operating expenses:
Selling expenses 67200
Administrative expenses 29500
Total operating expenses 96700
Net income 49100
b
Digital Vibe Manufacturing Company
Inventory Balances
For the Month Ended January 31
Inventory balances on January 31:
Materials 12600 =54600-42000
Work in process 46400 =42000+62800+88500-146900
Finished goods 30100 =146900-116800